Water flowing through a 1.7-cm-diameter pipe can fill a 300 L bathtub in 5.9 min.

What is the speed of the water in the pipe?

Express your answer in meters per second.
Answer #1

cross section area of pipe = (pi/4) x ( 1.7 /100)^2 = 2.26 x 10^-4 m^2

now
volume = 300 x 0.001 = 0.3 m^3 and
time = 5.9 x 60 = 354 sec
now
volume / time = cross section area x velocity
0.3 / 354= 2.26 x 10^-4 x v
v = 3.74 m/s